OpenStudy (anonymous):

How do i solve x+y=8 and y-x=2 so i can graph them?

OpenStudy (saifoo.khan):

they are perp lines.

OpenStudy (anonymous):

y = -x +8 and y = x+2

OpenStudy (anonymous):

you can plot them after that

OpenStudy (anonymous):

if im correct my paper is asking for the x and y intercept

OpenStudy (anonymous):

\[x+y=8\] x intercept is (8,0) and y intercept is (0,8)

OpenStudy (anonymous):

\[x-y=2\] x intercept is (2,0) and y -intercept is (0,-2)

OpenStudy (anonymous):

now how do i come to that? iwant to learn how to do it

OpenStudy (anonymous):

The x intercept is where the x value is 0, so plug 0 in for x and solve for y And vice versa for the y intercept